# Heterogeneity in Entity Matching: A Survey and Experimental Analysis

### v1

@@ 5.1.2 Entity Matching Models Evaluated
We evaluate four EM models in our experiments: DeepMatcher, DITTO, EM Transformer, and HierGAT. This selection balances practical considerations and architectural diversity. While earlier sections review a broad landscape of EM methods, pilot experiments showed that many recent models exhibit similar performance trends. Including all of them would add complexity without significantly altering conclusions. Thus, we focus on models that are actively maintained, run on modern libraries, and install without legacy dependencies—criteria that many older systems no longer meet.
The selected models span distinct design paradigms. DeepMatcher [50] is a widely-used baseline with a relatively simple architecture that continues to perform well, especially when sufficient labeled data is available. It combines hybrid attention over tokenized inputs, pre-trained word embeddings, and optional metadata to compute similarity scores. The remaining three methods—DITTO, EM Transformer, and HierGAT —introduce architectural innovations aimed at better handling heterogeneous EM (HEM). DITTO leverages input augmentation and a Transformer encoder, EM Transformer blends rule-based and learned matching strategies, and HierGAT uses graph-based contextual modeling. Together, these models offer a diverse and representative set of approaches for evaluating robustness under data heterogeneity.
@@ 5.2.1 Language & Terminology Heterogeneity with Synonyms
In the Company dataset (Figure 2(b)), initial AUC scores are slightly higher, and the performance drop from synonym injection is more gradual. HierGAT performs the worst in this setting, while DeepMatcher and DITTO lead. Figure 2(e) suggests that DeepMatcher benefits from the dataset’s well-structured attributes (e.g., “name”, “address”), where static embeddings suffice. DITTO remains strong due to its semantic generalization, whereas HierGAT’s attention mechanisms are less useful in strictly structured data.
@@ 5.2.3 Granularity & Resolution Heterogeneity with Hierarchical Distortion
The contrast with Figure 5 is notable. When the test data is distorted, critical features used for matching are obfuscated, leading to significant performance drops—often to the point where even human annotators would struggle. In contrast, when only the training data is distorted, robust models can still infer which attributes are most informative and learn effective representations. This allows them to generalize well to clean test data.
@@ 5.2.4 Heterogeneity Caused by Data Quality Differences
To assess sensitivity to noisy labels, we flipped a fraction of labels in the training set and measured performance on clean test data (Figure 8). As label noise increases, AUC declines across all models, but to varying degrees. In Fodor-Zagat, performance remains stable up to 10% noise before dropping sharply—especially for EM Transformer and HierGAT. In Walmart-Amz, all models degrade quickly, with DeepMatcher most affected. On iTunes-Amz, DITTO and HierGAT show stronger resilience, whereas EM Transformer and DeepMatcher degrade rapidly.
@@ 5.3 Key Findings, Limitations, and Implications
First, all forms of semantic heterogeneity—including language and terminology differences, granularity mismatches, and data quality issues—pose substantial challenges to entity matching. While advanced models like DITTO and HierGAT demonstrate greater resilience due to their use of contextualized embeddings and attention mechanisms, simpler architectures like DeepMatcher are highly sensitive to such variations, suffering steep performance declines in noisy or semantically inconsistent settings. This underscores the importance of using models that can capture deeper semantic and structural relationships.
@@
Third, model performance varies significantly by dataset. Complex or noisy datasets such as Walmart-Amz induce larger performance drops than simpler ones like iTunes-Amz. Tailoring methods to the characteristics of the data—e.g., attribute richness, schema complexity, or error patterns—can improve outcomes and guide model selection.
@@ 6 Conclusion and Future Research
Several key directions can guide future work on HEM:
  • EM in Data Lakes. Data lakes present a unique challenge for HEM due to their scale, diversity, and lack of predefined schema. Future research should address schema discovery, metadata integration, and multimodal entity matching by developing scalable algorithms that operate across unstructured and semi-structured formats. Methods must also accommodate evolving and hierarchical attributes while leveraging adaptive indexing and parallelism for large-scale deployments.
  • LLMs. LLMs offer strong potential for handling heterogeneity, particularly semantic variation. They enable few-shot or zero-shot EM and support alignment across modalities. However, challenges remain. Effective prompt engineering, robustness to domain shifts, and stable fine-tuning in low-resource settings are all open research problems. Future work should explore multimodal LLMs and retrieval-augmented architectures for EM, as well as continual adaptation to evolving data.
  • Knowledge Graphs (KGs). KGs help resolve semantic and structural inconsistencies by providing external context and mapping synonymous or ambiguous terms to canonical representations. Integrating KGs with EM models can improve explainability and performance, particularly in multilingual and cross-domain settings. Dynamic or domain-specific KGs can further aid in adapting to temporal and contextual shifts in data.
  • Human-in-the-Loop (HITL) and Explainability. Automated methods alone often fail in cases of context-sensitive or ambiguous matches. HITL systems allow users to provide guidance and resolve difficult cases. Future systems should incorporate interactive feedback loops and explanation tools that help users understand model predictions—especially in domains like healthcare and finance where accuracy and interpretability are critical.
  • Privacy and Security. Heterogeneous data often spans sensitive domains. EM systems must therefore support privacy-preserving mechanisms, such as differential privacy adapted for mixed data types, federated learning for decentralized matching, and secure multi-party computation. These techniques must balance utility and protection across structured, unstructured, and multimodal datasets.
  • Fairness and Inclusivity. Recent work has begun to address fairness in EM, exploring how biases emerge when datasets differ in completeness or representation across subgroups [65, 48, 52, 18, 49]. These studies have introduced fairness-aware models and evaluation protocols, laying a foundation for equitable EM. Yet, gaps remain: most approaches focus on static bias mitigation, neglecting fairness under evolving data, cross-domain heterogeneity, and intersectional subgroups. Future work should create EM systems that continuously monitor and adapt fairness, use causal reasoning to trace bias origins, and leverage user feedback to balance equity with accuracy. Techniques such as adversarial debiasing, dynamic re-weighting, and knowledge-driven disambiguation can move fairness-aware EM beyond isolated benchmarks toward robust, real-world deployment.
  • Evaluation and Benchmarking. Progress in HEM requires robust, diverse benchmarks. Current datasets insufficiently capture real-world complexity. New benchmarks should reflect multimodal, cross-lingual, and domain-specific scenarios. Evaluation protocols must move beyond traditional metrics to include fairness, explainability, and scalability. Synthetic data generation can help simulate heterogeneity scenarios systematically for deeper analysis.
Advancing EM in heterogeneous environments will require methods that are semantically aware, scalable, fair, privacy-conscious, and explainable. Addressing these challenges is essential for building trustworthy EM systems that are effective across real-world applications.

### v2

@@ 5.1.2 Entity Matching Models Evaluated
We evaluate four EM models in our experiments: DeepMatcher, DITTO, EM Transformer, and HierGAT. This selection balances practical considerations and architectural diversity. While earlier sections review a broad landscape of EM methods, pilot experiments showed that many recent models exhibit similar performance trends. Including all of them would add complexity without significantly altering conclusions. Thus, we focus on models that are actively maintained, run on modern libraries, and install without legacy dependencies–criteria that many older systems no longer meet.
The selected models span distinct design paradigms. DeepMatcher [84] is a widely-used baseline with a relatively simple architecture that continues to perform well, especially when sufficient labeled data is available. It combines hybrid attention over tokenized inputs, pre-trained word embeddings, and optional metadata to compute similarity scores. The remaining three methods–DITTO, EM Transformer, and HierGAT–introduce architectural innovations aimed at better handling heterogeneous EM (HEM). DITTO leverages input augmentation and a Transformer encoder, EM Transformer blends rule-based and learned matching strategies, and HierGAT uses graph-based contextual modeling. Together, these models offer a diverse and representative set of approaches for evaluating robustness under data heterogeneity.
@@ 5.2 Experimental Results
Unless stated otherwise, we repeat each stochastic setting (e.g., different random seeds and, when relevant, different perturbations) and report mean ROC AUC with $\pm 1$ standard deviation (shaded bands; mean $\pm$ SD). Our goal in this section is to assess robustness trends under controlled heterogeneity, not to over-interpret small gaps between similar methods. Although significance tests for AUC differences are possible, applying them across many datasets and conditions would require heavy multiple-comparison correction and can highlight trivial effects. We therefore focus on run-to-run variability and treat differences within that variability as inconclusive.
@@ 5.2.1 Language & Terminology Heterogeneity with Synonyms
In the ataset (Figure LABEL:fig:company-syn), initial AUC scores are slightly higher, and the performance drop from synonym injection is more gradual. HierGAT performs the worst in this setting, while DeepMatcher and DITTO lead. Figure LABEL:fig:company-syn-c suggests that DeepMatcher benefits from the dataset’s well-structured attributes (e.g., “name”, “address”), where static embeddings suffice. DITTO remains strong due to its semantic generalization, whereas HierGAT’s attention mechanisms are less useful in strictly structured data.
@@ 5.2.3 Granularity & Resolution Heterogeneity with Hierarchical Distortion
The contrast with Figure 5 is notable. When the test data is distorted, critical features used for matching are obfuscated, leading to significant performance drops–often to the point where even human annotators would struggle. In contrast, when only the training data is distorted, robust models can still infer which attributes are most informative and learn effective representations. This allows them to generalize well to clean test data.
@@ 5.2.4 Heterogeneity Caused by Data Quality Differences
To assess sensitivity to noisy labels, we flipped a fraction of labels in the training set and measured performance on clean test data (Figure 8). As label noise increases, AUC declines across all models, but to varying degrees. In Fodor-Zagat, performance remains stable up to 10% noise before dropping sharply–especially for EM Transformer and HierGAT. In Walmart-Amz, all models degrade quickly, with DeepMatcher most affected. On iTunes-Amz, DITTO and HierGAT show stronger resilience, whereas EM Transformer and DeepMatcher degrade rapidly.
@@
5.2.5 Impact of Representation Heterogeneity
To simulate representation heterogeneity at the schema level, we start from the clean versions of Fodor-Zagat, iTunes-Amz, and Walmart-Amz and evaluate three representative neural matchers: DeepMatcher, HierMatch [97], and RoTom [77]. Each model is trained once on the original column order and then evaluated under two test-time conditions: (i) the original column order and (ii) a randomly shuffled order of attributes for each record pair. For every model–dataset–condition combination we repeat evaluation 20 times with different random seeds (and, for the shuffled case, different permutations), and report the mean ROC AUC and standard deviation in Table 3; the “No Shuffle” columns correspond to the original, unpermuted test schema.
[…]
Walmart-Amz HierMatch 94.34 $\pm$ 0.00 80.51 $\pm$ 14.09 -13.83
DeepMatcher 80.31 $\pm$ 0.91 72.60 $\pm$ 11.41 -7.71
RoTom 96.31 $\pm$ 0.30 94.88 $\pm$ 0.82 -1.43
iTunes-Amz HierMatch 93.32 $\pm$ 0.00 74.47 $\pm$ 19.37 -18.85
DeepMatcher 98.42 $\pm$ 0.40 81.18 $\pm$ 11.25 -17.24
RoTom 99.22 $\pm$ 0.24 95.32 $\pm$ 4.38 -3.90
Fodor-Zagat HierMatch 100.00 $\pm$ 0.00 79.98 $\pm$ 16.47 -20.02
DeepMatcher 99.92 $\pm$ 0.08 88.63 $\pm$ 11.73 -11.29
RoTom 99.99 $\pm$ 0.01 99.95 $\pm$ 0.07 -0.04
[…]
In Table 3, we observe that HierMatch, DeepMatcher, and RoTom all attain high ROC AUC when evaluated on the original column order, but that HierMatch and DeepMatcher suffer substantial drops once the attributes in the test records are randomly permuted. In addition, the standard deviations of AUC in the shuffled condition are large for these two models, indicating that their predictions are highly unstable across different permutations of the same records. RoTom, in contrast, retains almost all of its performance under column shuffling, with only minor decreases in AUC and consistently low variance, showing that it is effectively robust to this type of schema heterogeneity.
A plausible explanation for this behavior is that HierMatch and DeepMatcher are designed around a fixed attribute layout: they encode each tuple as a sequence of attribute representations whose positions are implicitly tied to particular fields, and they rely on RNN/attention layers and attribute-level parameters that are not permutation invariant. When the order of columns changes at test time, the model still interprets position $i$ as “the $i$-th training attribute”, so semantically mismatched features are compared and the learned decision boundary no longer aligns with the input. RoTom, on the other hand, linearizes records into text with explicit column-name markers and is trained together with data-augmentation operators (including column shuffling) on top of a pre-trained language model. As a result, the model learns to condition primarily on the column labels and textual content rather than their order, which naturally results the strong permutation robustness seen in the shuffled setting.
@@ 5.3 Key Findings, Limitations, and Implications
First, all forms of semantic heterogeneity–including language and terminology differences, granularity mismatches, and data quality issues–pose substantial challenges to entity matching. While advanced models like DITTO and HierGAT demonstrate greater resilience due to their use of contextualized embeddings and attention mechanisms, simpler architectures like DeepMatcher are highly sensitive to such variations, suffering steep performance declines in noisy or semantically inconsistent settings. This underscores the importance of using models that can capture deeper semantic and structural relationships.
@@
Third, model performance varies significantly by dataset. Complex or noisy datasets such as Walmart-Amz induce larger performance drops than simpler ones like iTunes-Amz. Tailoring methods to the characteristics of the data–e.g., attribute richness, schema complexity, or error patterns–can improve outcomes and guide model selection.
@@
Sixth, our error analysis reveals three architectural failure modes that help explain the performance gaps observed across Figures 29. In the synonym and synonym-vs-random experiments, DeepMatcher exhibits out-of-vocabulary failure: it relies on fixed GloVe/FastText vectors and maps many GPT-4–generated synonyms to generic unknown tokens, whereas transformer-based models with subword tokenization maintain a usable semantic signal. In the attribute-noise experiments, HierGAT suffers from graph-propagation of noise, since its message-passing layers spread corrupted attribute values to neighboring nodes, degrading representations more severely than sequence-based DITTO. Finally, in the missing-data experiments, DeepMatcher’s RNN-based attention is brittle under MCAR/MAR/MNAR because removing key tokens disrupts temporal dependencies, while DITTO’s self-attention can redistribute mass to remaining informative tokens (e.g., from a missing “Brand” to the “Title”), preserving stable performance even at high missingness rates.
@@ 6 Conclusion and Future Research
Several key directions can guide future work on HEM. Below we focus on areas that remain underexplored even after adding dedicated sections for LLMs, multimodal EM, and benchmarking in the main body of the paper.
  • EM in Data Lakes. Data lakes produce extreme representation and structural heterogeneity due to schema drift, sparse or unreliable metadata, and files spanning structured, semi-structured, and unstructured formats. Prior work on dataset discovery and ER in lakes [9, 12, 35] shows that mismatched or incomplete schemas make even simple alignment tasks difficult. Our experiments (Section 5) confirm that neural models degrade significantly under such schema and granularity shifts. Future research should develop adaptive matchers that combine schema inference, metadata enrichment, and multimodal content-based signals. Useful building blocks include table-understanding models such as TURL and TaPas [143, 48]. Promising directions include: (i) continual-learning EM models that update as lake schemas evolve; (ii) unified embeddings that reconcile structured, text, and image attributes; and (iii) pipelines that fuse metadata with content signals for robust matching at scale.
  • Human-in-the-Loop and Explainability. HITL EM has been explored for resolving difficult or ambiguous matches [3, 7], and recent studies on explainable ER [76, 33] show its relevance in practice. However, these systems rarely account for heterogeneity-driven errors such as context shifts or representation mismatches. Our experiments identify such cases as persistent failure modes. Future HEM research should combine uncertainty-aware active learning [100] with explanations tailored to our heterogeneity taxonomy—for example, highlighting when mismatches arise from missing attributes, conflicting context, or schema differences. HITL pipelines should also support incremental updates of match rules and embeddings as users provide feedback, enabling more robust and interactive EM.
  • Privacy and Security. Privacy-preserving linkage has a long history [125, 19], and federated or distributed EM techniques [56, 104] are gaining attention. However, most current systems assume consistent schemas and data types across parties. Heterogeneous schemas, mixed modalities, and evolving attributes create new privacy challenges not addressed by existing work. Future directions include designing DP-aware blocking and matching methods that work across heterogeneous attributes, building privacy-preserving multimodal embeddings, and developing secure multi-party protocols that handle schema drift. HEM can offer a structured way to reason about how different forms of heterogeneity interact with privacy risk.
  • Fairness and Inclusivity. Recent work on fairness in EM [105, 80, 86, 30, 81] has shown that real-world EM pipelines can amplify disparities across subgroups. Our empirical results indicate that heterogeneity—such as differing levels of attribute completeness or domain-specific terminology—intensifies these fairness issues. Future research should develop fairness metrics and mitigation strategies that explicitly account for semantic, contextual, and structural heterogeneity. Promising directions include causal analysis to trace how heterogeneous attributes propagate bias, dynamic re-weighting or adversarial debiasing to maintain fairness as data evolves, and schema-informed balancing that adjusts for subgroup-specific representation gaps.
  • Robustness to Temporal and Schema Drift. Temporal evolution creates new forms of heterogeneity even within a single source. Prior work on temporal ER [17] shows that entity relationships and attribute semantics can change substantially over time. Future work should design drift-aware EM pipelines that detect and localize semantic and schema changes, maintain cross-version attribute alignment, and update matchers via continual or online learning. The heterogeneity taxonomy offers a natural framework for identifying which aspects of drift, including semantic, contextual and structural drift, are most impactful and for guiding how systems should adapt.
Future advances in HEM require methods that explicitly handle the forms of heterogeneity outlined in our taxonomy and adapt as these conditions change. Such developments are essential for building EM systems that remain robust and reliable in real, evolving data ecosystems.
7 Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used ChatGPT (OpenAI) to improve the readability and language of the manuscript. After using this tool, the authors reviewed and edited the content as needed and took full responsibility for the content of the published article.